OpenStudy (anonymous):

The product of two consecutive negative integers is 1122. What are the numbers? The product of two consecutive negative integers is 1122. What are the numbers? @Mathematics

OpenStudy (bahrom7893):

x*(x-1)=1122

OpenStudy (campbell_st):

well the integers are - x and -(x+1) so there product is \[-x \times -(x +1) = 1122\] or \[x^2 - x - 1122 = 0\] solve the quadratic using the general quadratic formula \[x = \frac{1 \pm 67}{2}\] so x = - 33 the other integer is -(33 + 1) = -34 so your numbers are -33 and -34